[figure: 9.7 cm]

THE RIGGED CONFIGURATIONS IN THE BETHE ANSATZ METHOD

W. Babiaczyk Physics Department, A. Mickiewicz University, Umultowska 85, 61-614 Poznań, Poland

It is known that using the Bethe Ansatz method one can calculate the exact energies for the XXX model consisting of spins s=\frac{1}{2}. The rigged configurations are the combinatorial objects which parametrize the Bethe vectors. In this report the bijection between the solutions of Bethe Ansatz equations performed in terms of Takahashi numbers and the rigged configurations for the isotropic Heisenberg ferromagnet consisting of eight spins is presented.

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : W. Babiaczyk

Address for correspondence : Physics Department, Adam Mickiewicz University, Umultowska 85, 61-614 Poznań, Poland

Email address : iwo@hoth.amu.edu.pl